### v2

@@ 2 Background and Definitions
Because, given $x_{n}$, the value of $x_{n+1}$ can be computed exactly, this is a deterministic process, and the set of iterates, $x_{0},x_{1},x_{2},\dots=\{x_{n}\}_{n=0}^{\infty}$, is called the trajectory of the initial value $x_{0}$ under the dynamical system given by the function $f$. A natural question is: given an initial value, $x_{0}$, what is the long-term behavior of the sequence $x_{0},x_{1},x_{2},\dots$? If the sequence converges to a finite value, it must be a fixed point, i.e. a value $x^{*}$ that satisfies $x^{*}=f(x^{*})$. So one of the first questions a dynamicist asks when encountering a dynamical system is: where are the fixed points? Other questions of interest include determining the stability of fixed points - whether nearby trajectories converge toward, or away from, fixed points - and looking at the limiting behavior of non-fixed points.
In a stochastic dynamical system, we consider a sequence of random variables: $\{X_{n}\}_{n=0}^{\infty}$. Given a value for the random variable $X_{n}$ in the state space, the value of $X_{n+1}$ can not be determined exactly because the law of evolution is subject to some noise. We are then interested in the sequence of distributions of these random variables. In particular, we are interested in the existence of an invariant distribution: a distribution of $X_{n}$ such that $X_{n+1}$ has the same distribution. If such a distribution exists, we are also interested in its stability; that is, whether other distributions converge to this invariant distribution in some manner. We define the random logistic map and the notions of invariant measures and stability more precisely below.
@@
Let $\mathcal{B}(S)$ denote the Borel sigma-algebra on $S$. We consider the logistic Markov process
@@
Suppose we have a distribution, $\mu$, on the state space, $S$. When we apply the process (1) to subsets of $S$, we induce a new distribution, $P\mu$ on the state space as follows. If $A\subset S$, then we define
@@
We can iterate the operator $P$ a positive number $m$ times, which we denote by $P^{m}$: if $X_{0}$ has distribution $\mu$ then $X_{m}$ has distribution $P^{m}\mu$. This leads us the analog to stability in the stochastic context.
@@
In the deterministic case, we know that, if the iterates of $x_{0}$ under a continuous function, $f$, converge to a point, $x$, then $x$ must be a fixed point of $f$. This result follows from the continuity of $f$. We can get a similar result in the stochastic setting, but we first need to define a notion of continuity. Suppose we have a Markov process on $S$ with transition probabilities given by $p(x,dy)$.
@@
Using weak continuity we get the following result ([14], Theorem 11.1).
@@
    $\frac{1}{n_{k}}\sum_{m=0}^{n_{k}-1}p^{(m)}(x,dy)\hbox{ converges weakly to }\pi_{x}\hbox{ as }k\rightarrow\infty$
@@
    If the limit above exists for all $x\in S$ and is independent of $x$ then $\pi=\pi_{x}$ is the unique invariant probability for $p$.
We can state this in terms of the operator $P$ given in Equation (2). Suppose a process has the Feller property and we have a measure $\mu$ on $S$. If we have weak convergence in terms of the average to a measure, $\pi$:
@@
then $\pi$ is an invariant measure under the operator $P$, i.e. $P\pi(A)=\pi(A)$ for $A\in{\cal B}(S)$. Furthermore, if Condition 2 holds in Proposition 1, then this invariant measure is unique.
@@
A Markov process $\{X_{n}\}_{n=0}^{\infty}$ is Harris irreducible with reference measure $\nu$ (or $\nu$-irreducible for short) if, for every $x\in S$ and every $B\in\mathcal{B}(S)$ with $\nu(B)>0$, there is $n\geq 1$ such that $p^{n}(x,B)>0$.
In words: a Markov process is Harris irreducible if every element in the state space eventually gets mapped to any set of positive probability.
In the theorems that follow, we require that the process be Harris irreducible with respect to any $sigma$-finite measure: the particular measure is not important. Therefore, to make the language more concise, we sometimes just say “Harris irreducible” to mean “Harris irreducible with respect to some $\sigma$-finite measure.”
@@
The theorem applies to more general settings, but this more specific version is all we need here. Harris irreducibility can then be used to show that this invariant distribution is unique.
Theorem 2 ([14], Theorem 2.C9.3).
If a Markov process is Harris irreducible with respect to some $\sigma$-finite measure $\nu$ on $S$ and has an invariant probability $\mu^{*}$, then $\mu^{*}$ is the unique invariant probability and, for any probability measure $\mu$ on S:
@@
Thus, if a Markov process is Harris irreducible, we know that any measure converges to a unique invariant measure, $\mu^{*}$, under the operator, $P$ “in the sense of Césaro”, i.e. the average of the sequence of measures conerges to $\mu^{*}$. To get asymptotic stability of the invariant meausure without this average, we need one more condition: aperiodicity.
@@
A Markov process is said to be strongly aperiodic if there is a set $A_{0}\subset S$ with $\nu(A_{0})=1$ and an $\alpha>0$ such that for all $x\in A_{0}$, $A$ a measurable subset of $S$, one has
$p(x,A)\geq\alpha\nu(A)$
Theorem 3 ([14], Theorem C9.2).
If a Markov process is Harris irreducible with respect to some $\sigma$-finite measure $\nu$ on $S$ and strongly aperiodic with invariant measure $\mu^{*}$ then, for any $\mu\in\mathcal{B}(S)$,
$\lim_{n\rightarrow\infty}\delta(P^{n}\mu,\mu^{*})=0$
@@
where $\lambda$ is distributed according to a Bernoulli distribution, and takes two possible values, $\alpha$ and $\beta$ (where $\alpha<\beta$). Thus, the process switches randomly between two quadratic maps. Bhattacharya and Rao demonstrate that if $0\leq\alpha<\beta\leq 1$, then $\delta_{0}$, the Dirac measure supported on 0, is the unique invariant probability on [0,1]. If $1<\alpha<\beta\leq 2$ then there is a unique invariant probability on (0,1), and this probability is nonatomic. If $2<\alpha<\beta<1+\sqrt{5}$ and $\alpha\in[8/\beta(4-\beta),\beta)$, then there exists a unique invariant probability on (0,1) which is nonatomic and has its support contained in $[1/2,\lambda/4]$.
@@
Recall that the logistic family: $f_{\lambda}(x)=\lambda x(1-x)$ undergoes a period-doubling cascade as $\lambda$ increases from 1 to approximately 3.56995. For higher values of $\lambda$, the family exhibits stable periodic orbits of odd period, following the Sarkowski sequence, culminating with a stable period-3 orbit when $\lambda=1+\sqrt{8}$. When $\lambda$ is greater than this value the maps exhibit chaotic behavior on a Cantor set of parameter values. In this section, we are interested in randomly perturbing the values of $\lambda$ in this chaotic regime: $\lambda\in(1+\sqrt{8},4)$. The logistic family is well-studied in the literature, and more details on its dynamics can be found in, for example, [16]
In [17], Athreya and Dai consider the stochastic logistic map and demonstrate that the Markov process $\{X_{n}\}$ is Feller continuous, which guarantees the existence of the invariant distribution (but not its stability) by Theorem
1. They also show that if $E[\log\lambda_{1}]>0,E[\log(4-\lambda_{1})]<\infty$, and if the process is Harris irreducible with respect to some $sigma$-finite measure , then the probability distribution converges in the sense of Césaro to a unique probability distribution on (0,1). We shall consider the case when $\lambda\sim\mathcal{U}(3.87,4)$ to ensure that the maps are in the chaotic regime. We show that, in this case, the stochastic logistic map satisfies the conditions laid out in [17] for convergence of the distribution in the sense of Céaround.
Remark 1.
Let $\lambda\sim\mathcal{U}(3.87,4)$. Then we claim that $E[\log\lambda]>0$ and $E[\log(4-\lambda)]<\infty$. Notice that if $\lambda\sim\mathcal{U}(3.87,4)$, then $E(\lambda)=3.935>0$. Similarly,
[…]
$\approx$ $-12,$
Thus $E[\log(4-\lambda)]$ is finite.
This remark implies that if the process $\{X_{n}\}$ is Harris irreducible, then there exists a unique probability distribution on $(0,1)$ to which other probability distributions converge to in the sense of Cesaro.
The following results prove the Harris irreducibility and aperiodicity of the Markov process, which then gives the asymptotic stability of the invariant measure.
Lemma 1
and its generalization
Lemma 2 build toward Theorem 4, which gives the Harris irreducibility of the Markov process. Theorem 5 shows that the process is strongly aperiodic and, hence, has a unique asymptotically stable invariant measure.
@@
Let $\lambda_{i}\sim\mathcal{U}(3.87,4)$ for all $i\geq 1$, $\alpha$ a fixed element in $(3.87,4)$, and $x^{*}$ such that $x^{*}=\alpha x^{*}(1-x^{*})$. (Equivalently, $x^{*}=1-\frac{1}{\alpha}$.) Then there exists $\eta>0$ such that for all $x\in I=(x^{*}-\eta,x^{*}+\eta)$ and $A\subset I$ with $m(A)>0$, then
@@
That is, there is a neighborhood $I$ around the fixed point for which every $x\in I$ has a positive probability of visiting each subset of $I$ of positive
Lebesgue
measure.
@@
Note that while we take the distribution of $\lambda$ on $(3.87,4)$ to be uniform, the same result will hold if we replace the uniform distribution with a distribution that has non-vanishing absolutely continuous part
on this interval.
The next lemma generalizes Lemma
1
to the case in which the map has a stable periodic orbit of any period.
@@
As before, let $\lambda_{i}\sim\mathcal{U}(3.87,4)$ for all $i\geq 1$. Suppose $\alpha$ is a parameter value in $(3.87,4)$ with an $m$-periodic point, $p$, so that $f^{(m)}(p,\alpha)=p$. Then there is an $\eta>0$ such that, for any $x\in I=(p-\eta,p+\eta)$ and $A\subset I$ with $m(A)>0$, we have
@@
$43.87<\left({\frac{p-\eta_{1}}{p+\alpha\eta_{1}}}\right)\alpha<\left({\frac{p+\eta_{1}}{p-\alpha\eta_{1}}}\right)\alpha<$ 4
@@
$\lim_{\eta_{1}\rightarrow 0}\frac{p-\eta_{1}}{p+\alpha\eta_{1}}=\lim_{\eta_{1}\rightarrow 0}\frac{p+\eta_{1}}{p-\alpha\eta_{1}}=1$
Denote points in the periodic orbit of $p$ as $p_{j}=f^{(j)}(p,\alpha)$. Since $p$ has period $m$, $f(p_{m-1},\alpha)=\alpha p_{m-1}(1-p_{m-1})=p$. Since $f(\cdot,\cdot)$ is continuous in both variables:
@@
    there is an $\eta_{2}>0$ such that, if $|y-p_{m-1}|<\eta_{2}$ then
    $|f(y,\alpha)-p|<\eta_{1}\alpha\Leftrightarrow|y(1-y)-\frac{p}{\alpha}|<\eta_{1}\,;$
@@
    given $\eta_{2}$ as in the previous statement, $\exists\,\eta_{3}$ so that $|x-p|<\eta_{3}$ and $|\lambda_{i}-\alpha|<\eta_{3}$ for $1\leq i<m\Rightarrow|X_{m-1}-p_{m-1}|<\eta_{2}$ where $X_{j+1}=\lambda_{j+1}X_{j}(1-X_{j})$ and $X_{0}=x$.
In other words, if the random logistic map starts close to $p$, and if all of the parameter values are close to $\alpha$, then the random orbit returns close to $p$ after $m$ iterates.
@@
Remark 3.
Lemma 2 shows that, if there is a parameter $\alpha$ in the support of the parameter values for which the map has a stable periodic orbit, then the process $\{X_{n}\}$ is Harris irreducible on $S$ with reference measure $\phi(\cdot)=m(\cdot\cap I)$, where $m$ is Lebesgue measure.
@@
with $\{\lambda_{n}\}_{n\geq 1}$ being i.i.d., $\lambda_{i}\sim\mathcal{U}(3.87,4)$, and $X_{0}$ an independent $S$-valued r.v. Then $\{X_{n}\}$ is irreducible with respect to the Lebesgue measure $m$ on $S$.
@@
By a result of Graczyk and Swiatek [18], the set of all $\lambda$ in $(0,4)$ such that $f(\cdot,\lambda)$ has a stable periodic orbit is dense. Thus, there is a $\alpha\in(3.87,4)$ and $10<x^{*}<$ 1 such that $x^{*}$ is a stable periodic point for $f(x,\alpha)=\alpha x(1-x)$. Let $\delta$ be such that $(\alpha-\delta,\alpha+\delta)\subset(3.87,4)$.
@@
We will next show that for any $x\in(0,1)$, there is a positive probability that $x$ is mapped to $K$ in one iterate. In fact, we will show that $p(x,K)=Prob(X_{1}\in K|X_{0}=x)=1$. Fix $x\in(0,1)$ and let $K^{c}$ be the complement of $K$ in $[0,1]$. Since $\lambda$ is uniformly distributed on $[3.87,4]$ and $m(K^{c})=0$ we have
@@
We now have that, since $E(\lambda)>0$ and $E(\log(4-\lambda))<\infty$, and the process is Harris irreducible, by [17], the process $\{X_{i}\}$ has a unique invariant measure that is stable in the sense of Cesaro. We conclude this section by demonstrating that the process is strongly aperiodic, and thus the invariant distribution is asymptotically stable with respect to the total variation norm.
@@
Let $\lambda_{min}=3.87$, and let $A_{0}=(1-\frac{1}{\lambda_{min}},1-\frac{1}{4})$ (that is, $A_{0}$ is the set of fixed points for the deterministic logistic maps corresponding to the possible parameter $\lambda$ values). Let $L_{1}$ be the length of $A_{0}$ (that is, $L_{1}=\frac{1}{\lambda_{min}}-\frac{1}{4}$ ) and let $L_{2}$ be the length of interval $(\lambda_{min},4)=4-\lambda_{\min}$. Let $\nu$ be normalized Lebesgue measure on $A_{0}$. Then $A_{0}$, by definition, has measure 1, and for any measurable set $A$ :
@@
Fix $x\in A_{0}$ and $A\subset S$. We want to show that $p(x,A)\geq c\,\nu(A)$ for some positive constant, $c$ (Definition 5). We have:

### v3

@@ 2 Background and Definitions
Because, given $x_{n}$, the value of $x_{n+1}$ can be computed exactly, this is a deterministic process, and the set of iterates, $x_{0},x_{1},x_{2},\dots=\{x_{n}\}_{n=0}^{\infty}$, is called the trajectory of the initial value $x_{0}$ under the dynamical system given by the function $f$. A natural question is: given an initial value, $x_{0}$, what is the long-term behavior of the sequence $x_{0},x_{1},x_{2},\dots$? If the function $f$ is continuous, and the sequence converges to a finite value, $x^{*}$, then it must be a fixed point, i.e. $x^{*}=f(x^{*})$. So one of the first questions a dynamicist asks when encountering a dynamical system is: where are the fixed points? Other questions of interest include determining the stability of fixed points - whether nearby trajectories converge toward, or away from, fixed points - and looking at the limiting behavior of non-fixed points.
In a stochastic dynamical system, we consider a sequence of random variables: $\{X_{n}\}_{n=0}^{\infty}$. Given a value for the random variable $X_{n}$ in the state space, the value of $X_{n+1}$ cannot be determined exactly because the law of evolution is subject to some noise. We are then interested in the sequence of distributions of these random variables. In particular, we are interested in the existence of an invariant distribution: a distribution of $X_{n}$ such that $X_{n+1}$ has the same distribution. If such a distribution exists, we are also interested in its stability; that is, whether other distributions converge to this invariant distribution in some manner. We define the stochastic logistic map and the notions of invariant measures and their stability more precisely below.
@@
Let $\mathcal{B}(S)$ denote the Borel $\sigma$-algebra on $S$. We consider the logistic Markov process
@@
Suppose we have a distribution, $\mu$, on the state space, $S$, which gives the distribution of $X_{n}$. When we apply the process (1) to subsets of $S$, we induce a new distribution, $P\mu$ on the state space as follows. If $A\subset S$, then we define
@@
We can iterate the operator $P$ a positive number $m$ times, which we denote by $P^{m}$: if $X_{0}$ has distribution $\mu$ then $X_{m}$ has distribution $P^{m}\mu$. This leads us to the analog of stability in the stochastic context.
@@
In the deterministic case, we know that, if the iterates of $x_{0}$ under a continuous function, $f$, converge to a point, $x$, then $x$ must be a fixed point of $f$, and this result follows from the continuity of $f$. We can get a similar result in the stochastic setting, but we first need to define a notion of continuity. Suppose we have a Markov process on $S$ with transition probabilities given by $p(x,dy)$.
@@
Using weak continuity we get the following result ([14], Theorem 11.1 of Chapter 2).
@@
    $\frac{1}{n_{k}}\sum_{m=0}^{n_{k}-1}p^{m}(x,dy)\hbox{ converges weakly to }\pi_{x}\hbox{ as }k\rightarrow\infty$
@@
    If, for all $x\in S$, there is a sequence of integers $\{n_{i}\}$ such that the limit above exists and is independent of $x$ then $\pi=\pi_{x}$ is the unique invariant probability for $p$.
We can state this proposition in terms of the operator $P$ given in Equation (2). Suppose a process has the Feller property and we have a measure $\mu$ on $S$. If we have weak convergence in terms of the average to a measure, $\pi$:
@@
then $\pi$ is an invariant measure under the operator $P$, i.e. $P\pi(A)=\pi(A)$ for $A\in{\cal B}(S)$. Furthermore, if the condition in Item 2 of Proposition 1 holds, then this invariant measure is unique.
@@
  1. (i)
    A Markov process $\{X_{n}\}_{n=0}^{\infty}$ is irreducible with respect to a non-zero $\sigma$-finite reference measure $\nu$ (or $\nu$-irreducible for short) if, for every $x\in S$ and every $B\in\mathcal{B}(S)$ with $\nu(B)>0$, there is $n\geq 1$ such that $p^{n}(x,B)>0$.
  2. (ii)
    The Markov process is $\phi$-recurrent if, for each $x\in S$ and $A\in\mathcal{B}(S)$ such that $\phi(A)>0$, we have
    $P(\eta_{A}<\infty|X_{0}=x)=1,\quad\hbox{ where }\;\eta_{A}=\inf\{n\geq 1:X_{n}\in A\}$
  3. (iii)
    If the Markov process is both $\nu$-irreducible and $\phi$-recurrent, then it is Harris recurrent.
  4. (iv)
    If a Harris recurrent process satisfies
    $\sup\{E(\eta_{A_{0}}|X_{0}=x):x\in A_{0}\}<\infty$
    for some $\phi$-recurrent set $A_{0}$, then the process is said to be positive Harris recurrent.
Note: the set $A_{0}$ is $\phi$-recurrent if $\phi(A_{0})>0$ and it is recurrent, i.e. $P(\eta_{A_{0}}<\infty|X_{0}=x)=1$ $\forall x\in{\mathcal{B}}(S)$.
In other words: a Markov process is $\nu$-irreducible if there is a positive probability that every element in the state space eventually gets mapped to any set of positive $\nu$ measure. It is $\phi$-recurrent if, with probability 1, it visits every set of positive $\phi$-measure in a finite number of steps.
@@
The theorem applies to more general settings, but this more specific version is all we need here. Irreducibility can then be used to show that this invariant distribution is unique using the following theorem:
Theorem 2.
[[14], Theorem C9.3 in Chapter 2] If a Markov process is irreducible with respect to some $\sigma$-finite measure $\nu$ on $S$ and has an invariant probability $\mu^{*}$, then the process is positive Harris recurrent, $\mu^{*}$ is the unique invariant probability and, for any probability measure $\mu$ on S:
@@
Thus, if a Markov process is irreducible and has some invariant measure, we know that any measure converges to a unique invariant measure, $\mu^{*}$, under the operator, $P$ “in the sense of Césaro”, i.e. the average of the sequence of measures converges to $\mu^{*}$. We also get positive Harris recurrence from this result.
To get asymptotic stability of the invariant measure without this average, we need one more condition: aperiodicity.
@@
A Markov process is said to be strongly aperiodic with respect to a $\sigma$-finite probability measure $\phi$ on $S$ if there is a set $A_{0}\subset S$ with $\phi(A_{0})=1$ and an $\alpha>0$ such that one has
  1. 1.
    for all $x\in S$, $P(\eta_{A_{0}}<\infty|X_{0}=x)=1$, i.e. the process is $A_{0}$-recurrent;
  2. 2.
    for all $x\in A_{0}$, $A$ a measurable subset of $S$, $p(x,A)\geq\alpha\phi(A)$.
Let $\{X_{n}\}$ be a Markov process on a state space $S$, and ${\mathcal{B}}$ the Borel subsets of $S$.
Theorem 3 ([15], Corollary 6.7).
Suppose $\{X_{n}\}$ is a positive Harris recurrent and aperiodic Markov process with stationary distribution $\pi$. Then for any initial distribution, $\mu$, of $X_{0}$:
$\lim_{n\rightarrow\infty}\delta(P^{n}\mu,\pi)=0$
@@
We note that the measures used in the definitions of irreducibility and aperiodicity could be different. (This fact will be used later in the proof of Theorem 5.) To see that this is the case, consider the following. Aperiodicity with respect to the measure $\nu$ implies that there is an $\alpha>0$ such that
$p(x,B)\geq\alpha\nu(B)\,\Rightarrow\,p(x,B)>0\;\hbox{ if }\nu(B)>0$
Thus, the process is $\nu$-irreducible. Therefore, in the statement of Theorem 3, we do not need to specify the measures used to show either irreducibility or aperiodicity.
@@
where $\lambda$ is distributed according to a Bernoulli distribution, and takes two possible values, $\alpha$ and $\beta$ (where $\alpha<\beta$). Thus, the process switches randomly between two quadratic maps. Bhattacharya and Rao demonstrate that if $0\leq\alpha<\beta\leq 1$, then $\delta_{0}$, the Dirac measure supported on 0, is the unique invariant probability on [0,1]. If $1<\alpha<\beta\leq 2$ then there is a unique invariant probability on (0,1), and this probability is nonatomic. If $2<\alpha<\beta<1+\sqrt{5}$ and $\alpha\in[8/\beta(4-\beta),\beta)$, then there exists a unique invariant probability on (0,1) which is nonatomic and has its support contained in $[1/2,(1+\sqrt{5})/4]$.
@@
Recall that the deterministic logistic family: $f(x,\lambda)=f_{\lambda}(x)=\lambda x(1-x)$ undergoes a period-doubling cascade as $\lambda$ increases from 1 to $\lambda_{2\infty}\approx$ 3.56995. Note that here iteration occurs only with respect to $x$; that is $f^{(m)}(x,\lambda)=f(f^{m-1}(x,\lambda),\lambda)$, where $\lambda$ remains fixed. For some higher values of $\lambda$, the family exhibits stable periodic orbits of odd period, following the Sarkowski sequence, culminating with a stable period-3 orbit when $\lambda=1+\sqrt{8}$. Interspersed between the values of $\lambda$ with stable periodic orbits are a set of large measure of parameter values for which there are no stable periodic orbits, and almost all initial conditions give rise to orbits with no finite period. Thus, $\lambda_{2\infty}$ marks the “onset of chaos”. As $\lambda$ increases beyond $1+\sqrt{8}$, the family of maps again experiences a period-doubling cascade, with stable orbits of period 6, 12, 24, 48, …, culminating at $\lambda_{3\infty}\approx 3.8495$.
When $\lambda$ is greater than this value, the maps exhibit chaotic behavior on a Cantor set of parameter values, with three “chaotic bands” of orbits merging into one chaotic band at $\lambda\approx 3.857$. In what follows, we are interested in randomly perturbing the values of $\lambda$ in this chaotic regime. To make sure we are well within the parameter range for which there is generically a single band chaotic attractor, we restrict $\lambda\in(3.87,4)$. The logistic family is well-studied in the literature, and more details on its dynamics can be found in, for example, [17]
In [18], Athreya and Dai consider the stochastic logistic map and demonstrate that the Markov process $\{X_{n}\}$ is Feller continuous, which guarantees the existence of an invariant distribution (but not its stability) by Theorem 1. They also give conditions for convergence of the distribution of the $X_{i}$ in the sense of Césaro which differ somewhat from the hypotheses of Theorem 2. Here we choose to follow Bhattacharya and Majumbdar ([14]) and Nummelin ([15]) by showing the irreducibility and aperiodicity of the Markov process in the chaotic regime. This, along with the existence of an invariant measure, gives the asymptotic stability of the invariant measure by Theorem 3.
Lemma 1 and its generalization Lemma 2 build toward Theorem 4, which shows that the Markov process is irreducible. Theorem 5 shows that the process is strongly aperiodic and thus, since it has an invariant measure by Theorem 1, the process has a unique asymptotically stable invariant measure.
@@
Let $\lambda\sim\mathcal{U}(3.87,4)$, $\alpha$ a fixed element in $(3.87,4)$, and $x^{*}\in(0,1)$ such that $x^{*}=\alpha x^{*}(1-x^{*})$. (Equivalently, $x^{*}=1-\frac{1}{\alpha}$.) Then there exists $\eta>0$ such that for all $x\in I=(x^{*}-\eta,x^{*}+\eta)$ and $A\subset I$ with $m(A)>0$, then
@@
That is, there is a neighborhood $I$ around the fixed point for which every $x\in I$ has a positive probability of visiting each subset of $I$ of positive Lebesgue measure.
@@
Note that while we take the distribution of $\lambda$ on $(3.87,4)$ to be uniform, the same result will hold if we replace the uniform distribution with a distribution that has non-vanishing absolutely continuous part on this interval.
The next lemma generalizes Lemma 1 to the case in which the map has a stable periodic orbit of any period.
@@
As before, let $\lambda_{i}\sim\mathcal{U}(3.87,4)$ for all $i\geq 1$. Suppose $\alpha$ is a parameter value in $(3.87,4)$ with an $M$-periodic point, $x^{*}$, so that $f^{(M)}(x^{*},\alpha)=x^{*}$. Then there is an $\eta>0$ such that, for any $x\in I=(x^{*}-\eta,x^{*}+\eta)$ and $A\subset I$ with $m(A)>0$, we have
@@
$43.87<\left({\frac{x^{*}-\eta_{1}}{x^{*}+\alpha\eta_{1}}}\right)\alpha<\left({\frac{x^{*}+\eta_{1}}{x^{*}-\alpha\eta_{1}}}\right)\alpha<$ 4
@@
$\lim_{\eta_{1}\rightarrow 0}\frac{x^{*}-\eta_{1}}{x^{*}+\alpha\eta_{1}}=\lim_{\eta_{1}\rightarrow 0}\frac{x^{*}+\eta_{1}}{x^{*}-\alpha\eta_{1}}=1$
Denote points in the periodic orbit of $x^{*}$ as $x^{*}_{j}=f^{(j)}(x^{*},\alpha)$. Since $x^{*}$ has period $M$, $f(x^{*}_{M-1},\alpha)=\alpha x^{*}_{M-1}(1-x^{*}_{M-1})=x^{*}$. Since $f(\cdot,\cdot)$ is continuous in both variables:
@@
    there is an $\eta_{2}>0$ such that, if $|y-x^{*}_{M-1}|<\eta_{2}$ then
    $|f(y,\alpha)-x^{*}|<\eta_{1}\alpha\Leftrightarrow\left|y(1-y)-\frac{x^{*}}{\alpha}\right|<\eta_{1}\,;$
@@
    given $\eta_{2}$ as in the previous statement, $\exists\,\eta_{3}$ so that $|x-x^{*}|<\eta_{3}$ and $|\lambda_{i}-\alpha|<\eta_{3}$ for $1\leq i<M\Rightarrow|X_{M-1}-x^{*}_{M-1}|<\eta_{2}$ where $X_{j+1}=\lambda_{j+1}X_{j}(1-X_{j})$ and $X_{0}=x$.
In other words, if the random logistic map starts close to $x^{*}$, and if all of the parameter values are close to $\alpha$, then the random orbit returns close to $x^{*}$ after $M$ iterates.
@@
with $\{\lambda_{n}\}_{n\geq 1}$ being i.i.d., $\lambda_{i}\sim\mathcal{U}(3.87,4)$, and $X_{0}$ an independent $S$-valued r.v. Then $\{X_{n}\}$ is irreducible with respect to the Lebesgue measure $m$ on $S$ conditioned on $(x^{*}-\eta,x^{*}+\eta)$.
@@
By a result of Graczyk and Swiatek [19], the set of all $\lambda$ in $(0,4)$ such that $f(\cdot,\lambda)$ has a stable periodic orbit is dense. Thus, there is an $\alpha\in(3.87,4)$ and $10<x^{*}<$ 1 such that $x^{*}$ is a stable periodic point for $f(x,\alpha)=\alpha x(1-x)$.
@@
We will next show that for any $x\in(0,1)$, there is a positive probability that $x$ is mapped to $K$ in one iterate. In fact, we will show that $p(x,K)=Prob(X_{1}\in K|X_{0}=x)=1$. Fix $x\in(0,1)$ and let $K^{c}$ be the complement of $K$ in $[0,1]$. Since $\lambda$ is uniformly distributed on $(3.87,4)$ and $m(K^{c})=0$ we have
@@
Using Theorems 1 and 2, we now have that, since the process is Feller continuous by [18], the process $\{X_{i}\}$ has a unique invariant measure that is stable in the sense of Césaro. We conclude this section by demonstrating that the process is strongly aperiodic, and thus the invariant distribution is asymptotically stable with respect to the total variation norm.
@@
Let $\lambda_{min}=3.87$, and let $A_{0}=(1-\frac{1}{\lambda_{min}},1-\frac{1}{4})$ (that is, $A_{0}$ is the set of fixed points for the deterministic logistic maps corresponding to the possible parameter $\lambda$ values). Let $L_{1}$ be the length of $A_{0}$ (that is, $L_{1}=\frac{1}{\lambda_{min}}-\frac{1}{4}$ ) and let $L_{2}$ be the length of interval $(\lambda_{min},4)$ (that is, $L_{2}=4-\lambda_{\min}$). Let $\nu$ be normalized Lebesgue measure on $A_{0}$. Then $A_{0}$, by definition, has measure 1, and for any measurable set $A$ :
@@
Fix $x\in A_{0}$ and $A\subset S$. We want to show that $p(x,A)\geq c\,\nu(A)$ for some positive constant, $c$ (Definition 5). Since $\nu(A)=\nu(A\cap A_{0})$ it suffices to demonstrate the above result for sets $A\subset A_{0}$.
Notice for any $A\subset A_{0}$ and with $\nu(A)>0$, and $x\in A_{0}$,
[…]
and
[…]
Therefore, for any $A\subset A_{0}$ and $x\in A_{0}$, $\frac{A}{x(1-x)}\subset(3.87,4).$
Given $x\in A_{0}$ and $A\subset A_{0}$, we therefore have: